OpenStudy (anonymous):

Which point lies on the line described by the equation below? y + 3 = 2(x - 1) A. (0, 0) B. (1, -4) C. (-1, -7) D. (2, 9) E. (1, -3) F. (2, 1)

OpenStudy (anonymous):

y+3=2x-2 put 0 for y and that's going be 3=2x-2 3=2x-2 +2 5=2x x=2 now put 0 for x y+3=2 2-3 y=1 (2,1)